Find the price for good Z and the quantity supply for good X (show all the calculations) if:


(i) The elasticity of supply is equal to 1 and the price increases from $40 to $50.


(ii) If the elasticity of demand is 0.5 and the quantity demanded decreases from 95.000 to 85.000.


(iii) Draw the graph and indicate the equilibrium price and quantity.


Price Per Tonne ($) Quantity Demanded Quantity Supplied

40 150 80

50 120 X

60 110 110

80 95 115

Z 85 120

110 80 140





Explain each breakdown please and watch out as some of the numbers have decimals. Thank you for the help in advance


1
Expert's answer
2022-02-09T18:58:43-0500

Demand and Supply

Qn. i

"Elasticity\ of\ Supply(e_s)=1,\ P(\$40\to\$50)"

"e_s=\frac{\% \Delta Q_s}{\% \Delta P}=1"

"{\% \Delta P}=\frac{50-40}{100}=0.1"

"1=\frac{\% \Delta Q_s}{0.1}"

"\% \Delta Q_s=\frac{X-80}{100}=0.1"

"\bold{Q_s (X)=90}"


Qn. ii

"Elasticity\ of\ Demand(e_d)=0.5,\ Q(85\to95)"

"e_d=\frac{\% \Delta Q_d}{\% \Delta P}=0.5"

"{\% \Delta Q_d}=\frac{85-95}{100}=-0.1"

"0.5=\frac{-0.1}{\% \Delta P}"

"\% \Delta P=\frac{-0.1}{0.5}=-0.2"

"\therefore\ P(Z)=\frac{Z-80}{100}=-0.2"

"Z-80=-20"

"\bold{Z=60}"


Qn. iii